# Communication Optimization for Distributed Training: Architecture, Advances, and Opportunities

### v1

@@ V Conclusion
Due to the dramatic increase in the number of parameters of large-scale deep neural network models, it is imperative to develop effective communication optimization methods in distributed training. In this article, we first provide a general distributed training architecture and the current three-layer paradigm of communication optimization consisting of parallelization strategy, CCL, and network. We then review the latest advances in the current three-layer paradigm. Furthermore, we analyze the shortcomings of the current paradigm and advocate a communication-efficient five-layer paradigm, highlighting great opportunities for “cross-layer” co-design. We hope to shed some light on future research on communication optimization for distributed training.

### v2

@@ V Conclusion
Due to the dramatic increase in the number of parameters of large-scale deep neural network models, it is imperative to develop effective communication optimization methods in distributed training. In this article, we first introduce a general distributed training architecture from the communication perspective and summarize the current three-layer paradigm of communication optimization, including parallelization strategy, CCL, and network. We review the latest advances in the current three-layer paradigm and summarize their important features in a table. Then, we look forward to future research directions from the perspective of collaborative design. We advocate a communication-efficient five-layer paradigm and collaborative design for heterogeneous resources. Finally, we propose some open issues. We believe that this article can shed some light on future research on communication optimization for distributed training.
Acknowledgments
This work was supported by NSFC Project under Grant 62132009 and Grant 62221003.